# Cross-Level Sensor Fusion with Object Lists via Transformer for 3D Object Detection

### v1

@@ Abstract
In automotive sensor fusion systems, smart sensors and Vehicle-to-Everything (V2X) modules are commonly utilized. Sensor data from these systems are typically available only as processed object lists rather than raw sensor data from traditional sensors. Instead of processing other raw data separately and then fusing them at object level, we propose an end-to-end cross-level fusion concept with Transformer, which integrates highly abstract object list information with raw camera images for 3D object detection. Object lists are fed into a Transformer as denoising queries and propagated together with learnable queries through the latter feature aggregation process. Additionally, a deformable Gaussian mask, derived from the positional and size dimensional priors from the object lists, is explicitly integrated into the Transformer decoder. This directs attention toward the target area of interest and accelerates model training convergence. Furthermore, as there is no public dataset containing object lists as a standalone modality, we propose an approach to generate pseudo object lists from ground-truth bounding boxes by simulating state noise and false positives and negatives. As the first work to conduct cross-level fusion, our approach shows substantial performance improvements over the vision-based baseline on the nuScenes dataset. It demonstrates its generalization capability over diverse noise levels of simulated object lists and real detectors.
@@ I Introduction
3D object detection is fundamental in autonomous driving and robotics. While vision-based methods have advanced, sensor fusion remains crucial for robust perception in dynamic scenarios by combining the strengths of multiple sensor modalities. Multi-sensor fusion systems are commonly categorized into three primary architectures based on different levels of abstraction: low-level, intermediate-level, and high-level fusion [castanedo2013fusion], respectively fusing raw sensor signals, extracted features, and decisions or object lists. Low-level fusion theoretically offers optimal state estimation [thomopoulos1987optimal, chen2003limits, chong2011fusion], where recent advancements in deep neural networks (DNNs), particularly Transformers, have shown transformative capability.
V2X technology is widely involved in collective perception for autonomous driving [ambrosin2019design], where the demand for data transmission efficiency has constrained the usability of low-level fusion. Furthermore, smart sensors [spencer2004smart] process data internally, such as MobileEye vision systems and Continental radars, and output only perceived object states other than raw data [matzka2008multifusion]. In these situations, fusion is only conducted at a high level [chong2000track]. No previous research has explored using end-to-end learning-based methods to fuse object lists with raw sensor data.
@@
We define object lists as structured sets of object detections containing position $(x,y,z)$, size $(l,w,h)$, orientation $(yaw)$, velocity $(v_{x},v_{y})$, and class information with estimation errors. The cross-level fusion facilitates object detection by leveraging prior information from object lists. Consequently, we validate its performance using the nuScenes dataset [nuscenes2019]. Our approach significantly outperforms the vision-based baseline, revealing the substantial advantages of early integration of highly abstract information with raw sensor data.
@@ II-A Object-level Fusion and Data-level Fusion
In object-level fusion (OF) distributed object detectors send the source object lists (object state estimation with error covariance) to a central fuser for more accurate decisions [castanedo2013fusion], as shown in Fig. 1. OF often deploys explicit data association and Bayesian methods, like the Kalman filter with diverse motion models.
OF enhances redundancy by incorporating multiple sensor sources, ensuring system reliability when facing individual sensor failure. Besides, it improves computational efficiency by reducing the data volume processed, as the fusion operates on structured object detections rather than raw data. Furthermore, object fusion facilitates modular scalability, allowing for the seamless integration of additional sensors into the system without significant architectural modifications [MathWorksFusion, matzka2008multifusion]. However, OF often faces data association errors due to limited object information, causing object state estimation inaccuracies.
In contrast, data-level fusion fully leverages detailed low-level data [MathWorksFusion]. Typically implemented using deep learning models, it follows three main approaches: early fusion, where raw sensor data is unified; late fusion, where data is processed separately before merging decisions; and intermediate fusion, which integrates features at an intermediate stage, balancing data interaction with computational efficiency [huang2022multi]. However, these methods are designed for dense and reach positional data like point clouds and are not directly applicable to object lists, which contain sparse, explicit information, such as bounding boxes. To effectively fuse object lists, alternative strategies are needed to harness their structured, concise nature.
@@ II-B Two-stage Detection and Proposal-level Fusion
The two-stage detection process typically involves two main steps: region proposal and region classification [girshick2014rich]. Initially, the system processes sensor data to produce feature maps, which are used to identify regions of interest (RoIs) by projecting 3D bounding boxes onto 2D feature maps. These methods use the object hypotheses generated in the initial detection stage to refine and validate object positions, enhancing the overall accuracy and robustness of the detection system. Notable works like MV3D [chen2017multi] and AVOD [ku2018joint] deploy multi-view aggregation for detection, while other methods incorporate Transformer decoders for advanced fusion. FUTR3D [chen2023futr3d] emphasizes a unified query structure called Modality-Agnostic Feature Sampler across different modalities and incorporates a Transformer decoder to avoid late fusion heuristics. TransFusion [bai2022transfusion] fuses LiDAR and image data, with the first decoder layer predicting bounding boxes from LiDAR and the second adaptively integrating image features. An image-guided query initialization strategy further enhances the detection of objects that are challenging to identify in point clouds. FusionRCNN [xu2023fusionrcnn] utilizes RoIPooling to obtain data samples within proposals and then employs intra-modality self-attention to enhance domain-specific features and a cross-attention mechanism to align both modalities. However, the region proposals and object queries, usually of a significant amount than the ground truth, still originate from raw data or features, which inherently contain substantially richer information and redundancy compared to the object lists.
@@ III Methodology
The main aim is to implement a cross-level sensor fusion approach to fuse object list information with camera data end-to-end. For the overall architecture as shown in Fig. 2, we retained the main structure of our baseline PETRv2 [liu2023petrv2], due to its query-based design and temporal modeling through a 3D position-aware attention mechanism. Images $I=\{I_{i}\in\mathbb{R}^{3\times H_{I}\times W_{I}},i=1,2,\dots,N\}$ from N camera views are processed by a backbone network (e.g., VoVNetCP [Lee_2020_CVPR]) to extract 2D multi-view features $F^{2d}=\{F_{i}^{2d}\in\mathbb{R}^{C\times H_{F}\times W_{F}},i=1,2,\dots,N\}$. 3D coordinate generator discretizes the camera frustum space into a 3D meshgrid, transforming these coordinates into the 3D world space using camera parameters. 2D features and 3D coordinates from neighboring frames are concatenated and processed by the 3D position encoder (PE) to generate 3D position-aware features. These features serve as keys and values in the Transformer decoder.
@@ III-A Object-centric Query Denoising
Object-centric query denoising is firstly introduced in DN-DETR [li2022dn] to tackle the issues with the instability of bipartite graph matching and inconsistent optimization goals during early training stages of DETR. This method initializes queries with noisy ground-truth bounding boxes, termed “denoising queries”, aiming to reconstruct the original, clean boxes. The mechanism of DN-DETR series aligns with our assumption for cross-level fusion of object lists with raw sensor data using Transformers. The object list can be seen as noisy samplings of the ground truth with with estimation errors, FPs, and FNs, which can be directly converted to DN queries for DETR-like detectors. Besides, the built-in data processing systems to produce the object lists are black box, providing no information to de-correlate the object lists for optimal fusion [matzka2008multifusion]. However, learning-based approach enable end-to-end training and optimization, without the need for manual feature engineering or understanding of the underlying data processing mechanisms. Denoising enables the model to learn more robust feature representations that are less sensitive to variations and perturbations and to focus its attention more effectively for more precise object positioning and classification. Meanwhile, the queries sourced from the object list are enriched with raw data feature during cross-attention, and prepared for concatenation with extracted object embedding to the data association module, while DN-DETR discards denoising queries during inference. Introducing multiple denoising groups provides diverse training scenarios with a larger variety of noise levels, which accelerates the convergence and also prevents the overfitting to the GT-like data.
Moreover, DN-DETR highlights that feeding ground truth-like information directly into the Transformer during training can cause over-reliance, degrading performance during inference. Hence, an extra attention mask regulating the visibility of two parts of queries is incorporated into the self-attention mechanism in the decoder to prevent information leakage [li2022dn]. In QDN, we discard the attention mask which prevents the self attention so that the detection queries can get useful information from the denoising part. The co-occurrence between raw data and object list facilitate the prediction of objects on raw data.
@@ III-B Explicit Attention Directing
In addition to implicitly learning through the attention mechanism, we utilize a deformable Gaussian mask to adjust positional attention during cross attention, inspired by Spatially Modulated Co-Attention (SMCA) [gao2021fast]. SMCA creates a soft link between the object list and the raw image data to dynamically emphasize relative position during the learning process.
While the SMCA approaches referenced in  [bai2022transfusion, gao2021fast] establish a one-to-one alignment between queries and targets constrained by a Gaussian mask around the initial query, PETRv2 generates a large amount of random queries without predefined links to the Gaussian circles around targets. To solve this issue, we implement a one-to-many assignment of detection queries to targets within the Gaussian mask. This allows queries to map to any target or even a no-object class after training.
Fig. 3 provides a detailed flowchart of the process to compute and integrate the Gaussian mask into the Transformer decoder. Initially, object position data from the object lists and position encoding of objects in the 2D images are ported to the mask generator. With the projection matrices of cameras, we project their corresponding positions in multi-view images, subsequently resulting in a 2D Gaussian weighting mask $M$ for each target. The mask generation process resembles the one applied in  [zhou2019objects]:
@@
Detection Backbone NDS $\uparrow$ mAP $\uparrow$ mATE $\downarrow$ mASE $\downarrow$ mAOE $\downarrow$ mAVE $\downarrow$ mAAE $\downarrow$ PETRv2 [liu2023petrv2] V2-99 0.4038 0.4136 0.7208 0.7094 1.5492 0.4084 0.1914 our model V2-99 0.5685 0.5499 0.6334 0.2558 0.5006 0.4860 0.1885
@@ IV-A Experimental Setup
We chose the nuScenes dataset [nuscenes2019] as our experimental data source, which is a popular benchmark widely used for 3D object detection and multi-object tracking tasks in large-scale autonomous driving. The nuScenes dataset provides synchronized data from 6 cameras surrounding the ego car, capturing 10 object categories for detection and 7 moving classes for tracking in a 360°-field of view at a frequency of 12Hz. It contains 700, 150 and 150 scenes respectively in the training, validation and test sets.
@@
Our model and data processing pipeline are based on the MMDetection3D [mmdet3d2020] framework, facilitating efficient code development. Besides the multi-view images, the 3D annotations are imported into our training/testing pipeline with MMDetection3D to generate pseudo-object lists. We manually adopted a unified setup for all training experiments regarding the object list generation with max noise standard deviation ratio of 0.06, max drop rate of 0.2, max false positive rate of 0.1, and max label change rate of 0.3.
@@ IV-B3 Robustness against Modality Loss
From an engineering perspective, when raw data is unavailable, the object list should be directly forwarded to the decision-making module. Therefore, we define modality loss here as the absence of object lists. QDN differs from DN by removing the attention mask regulating the visibility between DN-queries and randomly-initialized queries. The significant performance drop in detection, a decrease by 25% in mAP on our model under the absence of object lists, shows the reliability of Transformer decoder on the DN-queries during inference. Interestingly, the explicit prior integration via SMCA can slightly mitigate the degeneration. However, this strong sensitivity to presence of object lists is worthy of further exploration. A potential solution is to introduce masked-modality training as in CMT [yan2023cross], where object lists are randomly masked during training. While we have not yet implemented this, future work could investigate its effectiveness in improving robustness to missing object lists. Additionally, we investigate the direct application of the original denoising mechanisms for the fusion of object lists. However, the results were unexpectedly suboptimal. When incorporating auxiliary high-level modal information into DN-queries during inference, the performance degradation was even more pronounced than with QDN without object lists. When injecting denoised queries during inference, the learnable queries may deviate from their intended targets.
@@ IV-B4 Generalization with Real and Synthetic Object Lists
The findings in Tab. IV demonstrate our model’s generalization capacity across different object list sources. With synthetic object lists, performance correlates directly with noise levels, while maintaining robustness under moderate noise conditions. To validate our approach with realistic detections, we conducted experiments using object lists from actual 3D detectors. With CenterPoint [yin2021center] ( mAP 0.564, NDS 0.652), CLF achieves performance (mAP: 0.469, NDS: 0.531) comparable to moderate-noise synthetic data. PointPillars [lang2019pointpillars], with lower performance (mAP: 0.343, NDS: 0.491), still provides significant improvements on NDS over the vision-only baseline. These results confirm both the effectiveness of our cross-level fusion approach with real detector outputs and the fidelity of our POLG noise parameters as reasonable approximations of actual detector behavior.

### v2

@@ Abstract
In automotive sensor fusion systems, smart sensors and Vehicle-to-Everything (V2X) modules are commonly utilized. Sensor data from these systems are typically available only as processed object lists rather than raw sensor data from traditional sensors. Instead of processing other raw data separately and then fusing them at the object level, we propose an end-to-end cross-level fusion concept with Transformer, which integrates highly abstract object list information with raw camera images for 3D object detection. Object lists are fed into a Transformer as denoising queries and propagated together with learnable queries through the latter feature aggregation process. Additionally, a deformable Gaussian mask, derived from the positional and size dimensional priors from the object lists, is explicitly integrated into the Transformer decoder. This directs attention toward the target area of interest and accelerates model training convergence. Furthermore, as there is no public dataset containing object lists as a standalone modality, we propose an approach to generate pseudo object lists from ground-truth bounding boxes by simulating state noise and false positives and negatives. As the first work to conduct cross-level fusion, our approach shows substantial performance improvements over the vision-based baseline on the nuScenes dataset. It demonstrates its generalization capability over diverse noise levels of simulated object lists and real detectors.
@@ I Introduction
3D object detection is fundamental in autonomous driving and robotics. While vision-based methods have advanced, sensor fusion remains crucial for robust perception in dynamic scenarios by combining the strengths of multiple sensor modalities. Multi-sensor fusion systems are commonly categorized into three primary architectures based on different levels of abstraction: low-level, intermediate-level, and high-level fusion [4], respectively fusing raw sensor signals, extracted features, and decisions or object lists. Low-level fusion theoretically offers optimal state estimation [22, 5, 10], where recent advancements in deep neural networks (DNNs), particularly Transformers, have shown transformative capability.
V2X technology is widely involved in collective perception for autonomous driving [1], where the demand for data transmission efficiency has constrained the usability of low-level fusion. Furthermore, smart sensors [21] process data internally, such as MobileEye vision systems and Continental radars, and output only perceived object states other than raw data [20]. In these situations, fusion is only conducted at a high level [9]. No previous research has explored using end-to-end learning-based methods to fuse object lists with raw sensor data.
@@
We define object lists as structured sets of object detections containing position $(x,y,z)$, size $(l,w,h)$, orientation $(yaw)$, velocity $(v_{x},v_{y})$, and class information with estimation errors. The cross-level fusion facilitates object detection by leveraging prior information from object lists. Consequently, we validate its performance using the nuScenes dataset [3]. Our approach significantly outperforms the vision-based baseline, revealing the substantial advantages of early integration of highly abstract information with raw sensor data.
@@ II-A Object-level Fusion and Data-level Fusion
In object-level fusion (OF) distributed object detectors send the source object lists (object state estimation with error covariance) to a central fuser for more accurate decisions [4], as shown in Fig. 1. OF often deploys explicit data association and Bayesian methods, like the Kalman filter with diverse motion models.
OF enhances redundancy by incorporating multiple sensor sources, ensuring system reliability when facing individual sensor failure. Besides, it improves computational efficiency by reducing the data volume processed, as the fusion operates on structured object detections rather than raw data. Furthermore, object fusion facilitates modular scalability, allowing for the seamless integration of additional sensors into the system without significant architectural modifications [8, 20]. However, OF often faces data association errors due to limited object information, causing object state estimation inaccuracies.
In contrast, data-level fusion fully leverages detailed low-level data [8]. Typically implemented using deep learning models, it follows three main approaches: early fusion, where raw sensor data is unified; late fusion, where data is processed separately before merging decisions; and intermediate fusion, which integrates features at an intermediate stage, balancing data interaction with computational efficiency [14]. However, these methods are designed for dense and reach positional data like point clouds and are not directly applicable to object lists, which contain sparse, explicit information, such as bounding boxes. To effectively fuse object lists, alternative strategies are needed to harness their structured, concise nature.
@@ II-B Two-stage Detection and Proposal-level Fusion
The two-stage detection process typically involves two main steps: region proposal and region classification [13]. Initially, the system processes sensor data to produce feature maps, which are used to identify regions of interest (RoIs) by projecting 3D bounding boxes onto 2D feature maps. These methods use the object hypotheses generated in the initial detection stage to refine and validate object positions, enhancing the overall accuracy and robustness of the detection system. Notable works like MV3D [6] and AVOD [15] deploy multi-view aggregation for detection, while other methods incorporate Transformer decoders for advanced fusion. FUTR3D [7] emphasizes a unified query structure called Modality-Agnostic Feature Sampler across different modalities and incorporates a Transformer decoder to avoid late fusion heuristics. TransFusion [2] fuses LiDAR and image data, with the first decoder layer predicting bounding boxes from LiDAR and the second adaptively integrating image features. An image-guided query initialization strategy further enhances the detection of objects that are challenging to identify in point clouds. FusionRCNN [23] utilizes RoIPooling to obtain data samples within proposals and then employs intra-modality self-attention to enhance domain-specific features and a cross-attention mechanism to align both modalities. However, the region proposals and object queries, usually of a significant amount than the ground truth, still originate from raw data or features, which inherently contain substantially richer information and redundancy compared to the object lists.
@@ III Methodology
The main aim is to implement a cross-level sensor fusion approach to fuse object list information with camera data end-to-end. For the overall architecture as shown in Fig. 2, we retained the main structure of our baseline PETRv2 [19], due to its query-based design and temporal modeling through a 3D position-aware attention mechanism. Images $I=\{I_{i}\in\mathbb{R}^{3\times H_{I}\times W_{I}},i=1,2,\dots,N\}$ from N camera views are processed by a backbone network (e.g., VoVNetCP [17]) to extract 2D multi-view features $F^{2d}=\{F_{i}^{2d}\in\mathbb{R}^{C\times H_{F}\times W_{F}},i=1,2,\dots,N\}$. 3D coordinate generator discretizes the camera frustum space into a 3D meshgrid, transforming these coordinates into the 3D world space using camera parameters. 2D features and 3D coordinates from neighboring frames are concatenated and processed by the 3D position encoder (PE) to generate 3D position-aware features. These features serve as keys and values in the Transformer decoder.
@@ III-A Object-centric Query Denoising
Object-centric query denoising is firstly introduced in DN-DETR [18] to tackle the issues with the instability of bipartite graph matching and inconsistent optimization goals during early training stages of DETR. This method initializes queries with noisy ground-truth bounding boxes, termed “denoising queries”, aiming to reconstruct the original, clean boxes. The mechanism of DN-DETR series aligns with our assumption for cross-level fusion of object lists with raw sensor data using Transformers. The object list can be seen as noisy samplings of the ground truth with with estimation errors, FPs, and FNs, which can be directly converted to DN queries for DETR-like detectors. Besides, the built-in data processing systems to produce the object lists are black box, providing no information to de-correlate the object lists for optimal fusion [20]. However, learning-based approach enable end-to-end training and optimization, without the need for manual feature engineering or understanding of the underlying data processing mechanisms. Denoising enables the model to learn more robust feature representations that are less sensitive to variations and perturbations and to focus its attention more effectively for more precise object positioning and classification. Meanwhile, the queries sourced from the object list are enriched with raw data feature during cross-attention, and prepared for concatenation with extracted object embedding to the data association module, while DN-DETR discards denoising queries during inference. Introducing multiple denoising groups provides diverse training scenarios with a larger variety of noise levels, which accelerates the convergence and also prevents the overfitting to the GT-like data.
Moreover, DN-DETR highlights that feeding ground truth-like information directly into the Transformer during training can cause over-reliance, degrading performance during inference. Hence, an extra attention mask regulating the visibility of two parts of queries is incorporated into the self-attention mechanism in the decoder to prevent information leakage [18]. In QDN, we discard the attention mask which prevents the self attention so that the detection queries can get useful information from the denoising part. The co-occurrence between raw data and object list facilitate the prediction of objects on raw data.
@@ III-B Explicit Attention Directing
In addition to implicitly learning through the attention mechanism, we utilize a deformable Gaussian mask to adjust positional attention during cross attention, inspired by Spatially Modulated Co-Attention (SMCA) [12]. SMCA creates a soft link between the object list and the raw image data to dynamically emphasize relative position during the learning process.
While the SMCA approaches referenced in  [2, 12] establish a one-to-one alignment between queries and targets constrained by a Gaussian mask around the initial query, PETRv2 generates a large amount of random queries without predefined links to the Gaussian circles around targets. To solve this issue, we implement a one-to-many assignment of detection queries to targets within the Gaussian mask. This allows queries to map to any target or even a no-object class after training.
Fig. 3 provides a detailed flowchart of the process to compute and integrate the Gaussian mask into the Transformer decoder. Initially, object position data from the object lists and position encoding of objects in the 2D images are ported to the mask generator. With the projection matrices of cameras, we project their corresponding positions in multi-view images, subsequently resulting in a 2D Gaussian weighting mask $M$ for each target. The mask generation process resembles the one applied in  [26]:
@@
Detection Backbone NDS $\uparrow$ mAP $\uparrow$ mATE $\downarrow$ mASE $\downarrow$ mAOE $\downarrow$ mAVE $\downarrow$ mAAE $\downarrow$ PETRv2 [19] V2-99 0.4038 0.4136 0.7208 0.7094 1.5492 0.4084 0.1914 our model V2-99 0.5685 0.5499 0.6334 0.2558 0.5006 0.4860 0.1885
@@ IV-A Experimental Setup
We chose the nuScenes dataset [3] as our experimental data source, which is a popular benchmark widely used for 3D object detection and multi-object tracking tasks in large-scale autonomous driving. The nuScenes dataset provides synchronized data from 6 cameras surrounding the ego car, capturing 10 object categories for detection and 7 moving classes for tracking in a 360°-field of view at a frequency of 12Hz. It contains 700, 150 and 150 scenes respectively in the training, validation and test sets.
@@
Our model and data processing pipeline are based on the MMDetection3D [11] framework, facilitating efficient code development. Besides the multi-view images, the 3D annotations are imported into our training/testing pipeline with MMDetection3D to generate pseudo-object lists. We manually adopted a unified setup for all training experiments regarding the object list generation with max noise standard deviation ratio of 0.06, max drop rate of 0.2, max false positive rate of 0.1, and max label change rate of 0.3.
@@ IV-B3 Robustness against Modality Loss
From an engineering perspective, when raw data is unavailable, the object list should be directly forwarded to the decision-making module. Therefore, we define modality loss here as the absence of object lists. QDN differs from DN by removing the attention mask regulating the visibility between DN-queries and randomly-initialized queries. The significant performance drop in detection, a decrease by 25% in mAP on our model under the absence of object lists, shows the reliability of Transformer decoder on the DN-queries during inference. Interestingly, the explicit prior integration via SMCA can slightly mitigate the degeneration. However, this strong sensitivity to presence of object lists is worthy of further exploration. A potential solution is to introduce masked-modality training as in CMT [24], where object lists are randomly masked during training. While we have not yet implemented this, future work could investigate its effectiveness in improving robustness to missing object lists. Additionally, we investigate the direct application of the original denoising mechanisms for the fusion of object lists. However, the results were unexpectedly suboptimal. When incorporating auxiliary high-level modal information into DN-queries during inference, the performance degradation was even more pronounced than with QDN without object lists. When injecting denoised queries during inference, the learnable queries may deviate from their intended targets.
@@ IV-B4 Generalization with Real and Synthetic Object Lists
The findings in Tab. IV demonstrate our model’s generalization capacity across different object list sources. With synthetic object lists, performance correlates directly with noise levels, while maintaining robustness under moderate noise conditions. To validate our approach with realistic detections, we conducted experiments using object lists from actual 3D detectors. With CenterPoint [25] ( mAP 0.564, NDS 0.652), CLF achieves performance (mAP: 0.469, NDS: 0.531) comparable to moderate-noise synthetic data. PointPillars [16], with lower performance (mAP: 0.343, NDS: 0.491), still provides significant improvements on NDS over the vision-only baseline. These results confirm both the effectiveness of our cross-level fusion approach with real detector outputs and the fidelity of our POLG noise parameters as reasonable approximations of actual detector behavior.